OpenStudy (anonymous):

Find the derivative of the function. y = (5x - 4)^4(8x^2 - 5)^-3

OpenStudy (anonymous):

use chain rule twice.

OpenStudy (mathteacher1729):

Also, product rule. :)

OpenStudy (mathteacher1729):

Or if you wanna go nuts, you can use the quotient rule as well: \[y = \frac{(5x-4)^4}{(8x^2-5)^3}\]

OpenStudy (anonymous):

do you understand that?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

chain rule: first, you take derivative of (5x-4)^4 = 4(5x-4)^3 multiply by derivative of (5x): 5 no change to second multiplier second step - first multiplier with no change * derivative of (8x^2-5)^(-3) : -3*(8x^2-5)^(-4) multiplied by derivative of (8x^2): 16x

OpenStudy (anonymous):

will give you: y'=4*5*(5x-4)^3 *(8x^2-5) -(5x-4)^4 *3*16x*(8x^2-5)^(-4) simplify

OpenStudy (anonymous):

try to do it yourself on paper & let me know

OpenStudy (anonymous):

ok